# Recent Advancements and Challenges of Turkic Central Asian Language Processing

### v2

@@ 1 Introduction
Turkic languages are spoken by around 200 million people globally, including over 60 million native speakers of Kazakh, Uzbek, Kyrgyz, and Turkmen. Due to their geographic, historical, and linguistic proximity, these languages face similar NLP challenges, ranging from limited quality data to restricted research opportunities.
@@
This paper aims to provide an overview of existing resources and suggest directions for future research to support both those utilizing current resources and those developing new ones. We also seek to highlight current resource needs, addressing which of those could be particularly crucial in advancing the Turkic Central Asian languages toward a higher-resource status.
@@ 2 Related Work
Recently, substantial efforts have been made to consolidate domain knowledge for Turkic languages via evaluation frameworks (Akın and Akın, 2007), linguistic analysis tools Abdurakhmonova et al. (2022), and NLP technology assessments Mirzakhalov et al. (2021); Maxutov et al. (2024). However, there is still a lack of comprehensive research summarizing the available data and language processing tools, especially for Central Asian Turkic languages. While state-of-the-art advancements in speech recognition and machine translation exist for some languages Bekarystankyzy et al. (2024); Yeshpanov et al. (2024b), no cross-linguistic comparisons have been conducted. A detailed survey could provide a valuable foundation for comparison and help define future research directions.
Figure 1: Distributions of numbers of Kazakh, Uzbek, Kyrgyz, and Turkmen native speakers among all Turkic language speakers. Numbers in the legend are approximates. Source: Wikipedia contributors 2024.
@@
Table 1: High-level overview of the Turkic Central Asian languages’ differences. Sources: Adali , Johanson .
@@ 3.1 Overview
Typology has a potential to improve language processing and transfer learning Ponti et al. (2019), and learning from similar languages is overall beneficial for the latter Zoph et al. (2016). Therefore, providing an overview of the linguistic features of Turkic languages may help identify potential similarities and challenges in their processing. While in depth linguistic comparison of each language with their richer counterpart, Turkish, lies beyond the scope of the paper, a basic summary is provided in the Table Table 1 to identify potential difficulties in the application of transfer learning from Turkish.
@@ 3.2 Similarities and Differences
Analyzing the linguistic differences in Table Table 1, such as variations in the number of pronouns or vowel sounds, one can infer that Kazakh and Kyrgyz are typologically closer to each other than Kazakh is to Uzbek, or Uzbek is to Kyrgyz; similarly, Turkmen is grammatically closest to Turkish. This suggests not only potential differences in transfer learning efficacy from Turkish, but also the possibility of successful transfer learning within the Central Asian group itself - likely making transfer between certain pairs, such as Kazakh and Kyrgyz, more effective than between others.
@@ 4 Datasets Availability
Open-source language resources enable researchers to scale and reuse data, reducing overall time spent on corpora collection and evaluation. However, data access varies from low to almost non-existent for the languages considered in the paper. A full summary of all the reviewed datasets and their primary features can be found in the Appendix A.
@@ 4.1 Sources of Data and Stakeholders
While developing a comprehensive dataset classification system would lie beyond the scope of the current paper, it is important to identify potential data collection sources and main stakeholders of the process. While a significant effort has been made by the researchers within the Central Asian countries themselves (e.g., most notable research on the Kazakh language is contributed by the Institute of Smart and Intelligent Systems (ISSAI) of Nazarbayev University), outside effort is driven primarily by the studies of NLP applications for Chinese minority languages Du et al. (2023) or Turkic languages in general Mirzakhalov et al. (2021). Additionally, while the researchers focus on collection of handcrafted high-quality datasets, other, potentially lower-quality data, is available via multilingual datasets crawled from the web, such as Common Crawl (CC) or OSCAR Ortiz Suárez et al. (2020). One more potential data source is the machine-translated content from high-resource languages; however, assessing the exact quantity of such data poses to be challenging, so the survey does not focus on it. Thus, there is a greater variety of data sources and stakeholders than might initially appear.
@@ 4.2 Kazakh Language Datasets
Unannotated Datasets and Corpora Collections. One of the largest available dataset of unannoted text articles in Kazakh is the Kazakh Language Corpus Makhambetov et al. (2013), containing a broad range of written text on a variety of topics. Additionally, annotated sub-corpora with linguistics and other language features are available within the same dataset. Another large collection of texts is the Kazakh National Language Corpus with over 22 thousand documents available.
Linguistic Features Datasets. Among the four surveyed languages, Kazakh has the most diverse and extensive datasets available. Besides collections of morphologically and syntactically-annotated data like Almaty Corpus of Kazakh Language and UD treebank Tyers and Washington (2015); Makazhanov et al. (2015), there exist also more task-specific text corpora.
@@ 4.3 Uzbek Language Datasets
Other Datasets. Some additional datasets, including multimodal and/or general unannotated data also exist. For example, one of the multimodal datasets is the open-source Uzbek Speech Corpus Musaev et al. (2021). On the other hand, there is the Uzbek corpus, which includes a large collection of educational, scientific, official, and artistic texts together with a morphological database and dictionaries, and an Uzbek Community corpus, collected by Leipzig University and containing over 660 thousand sentences of community data only.
@@ 4.4 Kyrgyz Language Datasets
Unannotated Datasets and Corpora Collections. One of the largest open-source datasets for Kyrgyz is the Manas-UdS corpus of over 84 literary texts in 5 genres, marked with lemmas and parts of speech Kasieva et al. (2020). For most other datasets, linking them directly to publications is not possible; however, some of them are available on Github. One of those is the KyrgyzNews dataset with over 250 thousand scraped news. Leipzig University has also compiled a corpora of over 3 million sentences from publicly available web sources.
Task-specific datasets. The few available task-specific datasets also can only be found on Github. One such example, for the task of NER, is the NER dataset that is currently under development by the researchers from Kyrgyz State Technical University (KSTU). Another example is the hand-written letters dataset (Kyrgyz MNIST equivalent) also available on Github.
@@ 4.5 Turkmen Language Datasets
The data landscape of Turkmen language is even more scarce. Besides the Leipzig University’s corpora of over 270 thousand sentences of web-scraped data, other resources are practically non-existent. With only a few dictionaries and poetry and literature collections collected by enthusiasts and available in an open-source fashion on Github, Turkmen language falls largely behind every other Turkic language of Central Asia in terms of data availability.
@@ 4.6 Web-Scraped Datasets
All the above-mentioned languages are also represented in multilingual datasets predominantly scraped from the web. In particular, all the languages are available in CC100 Wenzek et al. (2020), WikiAnn Pan et al. (2017), and OSCAR Ortiz Suárez et al. (2020) datasets. Most of other web-scraped datasets present online are either subsets of the bigger multilingual datasets (like CC100 or OSCAR), or are scrapes from newspapers and websites available on the Internet, with many of them not being open-source (for example, the kkWaC dataset with over 139 million Kazakh words). However, as noted by Kreutzer et al. (2022), such datasets should be used with caution, as the quality of low-resource language data in them may be significantly lower. With only a few hundred labeled instances, even a 5-10 percent rate of mislabeled or grammatically incorrect sentences can substantially impact model performance in these languages.
@@ 4.8 Parallel Corpora
Additionally, we would like to comment on the availability of parallel corpora for the languages studied. Apart from the Kazakh-Russian sign language corpora already mentioned, only a few parallel text datasets exist for Uzbek-Kazakh (Allaberdiev et al., 2024) and Kazakh-Russian (Kozhirbayev and Islamgozhayev, 2023) machine translation and ASR. However, parallel data is crucial for advancing one of the most prominent NLP technologies - translation tools.
@@ 4.9 Classifying Languages by Data Availability
The resources available for each language are far from abundant. If one was to categorize them according to the classification suggested by Joshi et al. (2020), Kazakh would most likely fit into “The Rising Star” category, with its substantial presence in the web and good variety of multimodal datasets. However, the research on this language is pushed back by a lack of labeled data for downstream tasks. Uzbek, given its recent developments and greater variability in terms of the linguistic resources available, would be categorized as “The Hopeful,” given that in the next years the efforts for collecting the datasets will not fade. Kyrgyz and Turkmen, unfortunately, not being sufficiently backed up by streamlined research efforts, would be classified as “The Scraping-Bys,” with the future of their data collection processes yet unclear.
@@ 5.2 Internet Access
Limited Internet access is the second reason for data scarcity in the region. As stated before, while a substantial amount of data has been gathered by NLP researchers from the Internet based sources, only 38 and 21 percent of users in Kyrgyzstan and Turkmenistan respectively have access to the global net. While the situation is somewhat better in Uzbekistan (with 55 percent of population being able to access the medium) and significantly better in Kazakhstan (around 79 percent of population), limited connectivity hinders users’ abilities to contribute to open-source encyclopedias, blogs, news sites, and more. Additionally, the lack of digitalized resources, such as electronic books, journals, audio transcripts, and video recordings, restricts information sharing within the region.
@@ 8.1 Kazakh Language Technologies
Linguistic Analysis. A variety of tools have been developed for linguistic analysis and normalization of texts in Kazakh Yessenbayev et al. (2020). This research paved the way for the first advances in the most prominent areas of NLP, like machine translation and ASR, that are well represented in Kazakh language.
@@ 8.2 Uzbek Language Technologies
ASR. While Uzbek does not enjoy the same variety of machine translation tools as Kazakh, presence of automatic speech recognition technology is somewhat comparable, with the most recent work contributed by Musaev et al. (2021), with their model performing at the level of 14.3 percent word error rate.
Fine-Tuning Existing Models. Uzbek also enjoys a better variety of pre-trained and fine-tuned models, including UzBERT Mansurov and Mansurov (2021), capable of outperforming mBERT by leveraging at least 11 times more of language specific data; UzRoberta, pre-trained on roughly 2 million news articles Davronov and Adilova (2024); BERTBek, a model, improving on the UzBERT decision by training on Latin script Kuriyozov et al. (2024), and a compact and fine-tuned variation of a mT5 model. In general, with Uzbek still catching up on data availability, there is potential for the language to soon enjoy a better variety of NLP technologies.
@@ 8.3 Kyrgyz and Turkmen Languages Technology
Unfortunately, the situation for Kyrgyz and Turkmen looks less promising. With little to no work in terms of machine translation or automatic speech recognition, Kyrgyz enjoys little variety of technology for text classification Alekseev et al. (2023) and NER fine-tuned on WikiAnn data. Turkmen offers even less language-specific technology, since it is has been mainly studied within the scope of comparative studies of Turkic languages. For example, some work on machine translation evaluation was done in the effort to build the infrastructure for Turkic languages, but nowadays this approach seems outdated (Alkim and Çebi, 2019). Overall, neither fine-tuning or pre-training of already available models and architectures have been researched for the majority of basic NLP tasks both in Kyrgyz and Turkmen.
@@ 9 Future Work Areas
Additionally, given the incorporation of LLMs in the field, potential of using them for data augmentation or annotation might happen to be another important direction of research in comparison with more expensive methods of human data collection and annotation.
@@ 11 Limitations
Appendix A Datasets

### v3

@@ 1 Introduction
Turkic languages are spoken by approximately 200 million people worldwide, with a significant concentration in Central Asia (see detailed breakdown of the number of speakers in Figure 1). While Turkish is the most resourceful language in the family, this paper focuses on less-resourced Turkic languages that are geographically, historically, and linguistically closer to one another in Central Asia. These languages represent an important subset of the Turkic family, spoken by approximately 80 million people in the region.
@@
This paper aims to provide an overview of existing resources and suggest directions for future research to support both those utilizing current resources and those developing new ones (for the search strategy details see Appendix A). We also seek to highlight current resource needs, addressing which of those could be particularly crucial in advancing the Turkic Central Asian languages toward a higher-resource status.
@@ 2 Related Work
Recently, substantial efforts have been made to consolidate domain knowledge for Turkic languages via linguistic analysis tools Akın and Akın (2007); Abdurakhmonova et al. (2022), and NLP technology assessments Mirzakhalov et al. (2021); Maxutov et al. (2024). However, there is still a lack of comprehensive research summarizing the available data and language processing tools, especially for Central Asian Turkic languages. While state-of-the-art advancements in speech recognition and machine translation exist for some languages Bekarystankyzy et al. (2024); Yeshpanov et al. (2024b), no cross-linguistic comparisons have been conducted. A detailed survey could provide a valuable foundation for comparison and help define future research directions.
@@
Table 1: High-level overview of the Turkic Central Asian languages’ differences. Sources: https://ecosystem.education/doc/Turkic%20Diller-$S$.pdf, https://www.britannica.com/topic/Turkic-languages
Figure 1: Distributions of numbers of Kazakh, Uzbek, Kyrgyz, and Turkmen native speakers among all Turkic language speakers. Numbers in the legend are approximates. Source: https://en.wikipedia.org/w/index.php?title=Languages_of_Asia&oldid=1230214231
@@ 3.1 Overview
Typology has a potential to improve language processing and transfer learning Ponti et al. (2019), and learning from similar languages is overall beneficial for the latter Zoph et al. (2016). Therefore, providing an overview of the linguistic features of Turkic languages may help identify potential similarities and challenges in their processing. While in depth linguistic comparison of each language with their richer counterpart, Turkish Çöltekin et al. (2022), lies beyond the scope of the paper, a basic summary is provided in Table 1.
@@ 3.2 Similarities and Differences
Analyzing the linguistic differences in Table 1, such as variations in the number of pronouns or vowel sounds, one can infer that Kazakh and Kyrgyz are typologically closer to each other than Kazakh is to Uzbek, or Uzbek is to Kyrgyz; similarly, Turkmen is grammatically closest to Turkish. This suggests not only potential differences in transfer learning efficacy from Turkish, but also the possibility of successful transfer learning within the Central Asian group itself - likely making transfer between certain pairs, such as Kazakh and Kyrgyz, more effective than between others.
@@ 4 Datasets Availability
Open-source language resources enable researchers to scale and reuse data, reducing overall time spent on corpora collection and evaluation. However, data access varies from low to almost non-existent for the languages considered in the paper. A full summary of all the reviewed datasets and their primary features can be found in the Appendix B.
@@ 4.1 Sources of Data and Stakeholders
While developing a comprehensive dataset classification system would lie beyond the scope of the current paper, it is important to identify potential data collection sources and main stakeholders of the process. While a significant effort has been made by the researchers within the Central Asian countries themselves (e.g., most notable research on the Kazakh language is contributed by the Institute of Smart and Intelligent Systems (ISSAI) of Nazarbayev University), outside effort is driven primarily by the studies of NLP applications for Chinese minority languages Du et al. (2023) or Turkic languages in general Mirzakhalov et al. (2021). Additionally, while the researchers focus on collection of handcrafted high-quality datasets, other, potentially lower-quality data, is available via multilingual datasets crawled from the web, such as Common Crawl (CC) [^1] https://www.commoncrawl.org or OSCAR Ortiz Suárez et al. (2020). One more potential data source is the machine-translated content from high-resource languages; however, assessing the exact quantity of such data poses to be challenging, so the survey does not focus on it. Thus, there is a greater variety of data sources and stakeholders than might initially appear.
@@ 4.2 Kazakh Language Datasets
Unannotated Datasets and Corpora Collections. One of the largest available dataset of unannoted text articles in Kazakh is the Kazakh Language Corpus Makhambetov et al. (2013), containing a broad range of written text on a variety of topics. Additionally, annotated sub-corpora with linguistics and other language features are available within the same dataset. Another large collection of texts is the Kazakh National Language Corpus [^2] https://qazcorpora.kz with over 22 thousand documents available.
Linguistic Features Datasets. Among the four surveyed languages, Kazakh has the most diverse and extensive datasets available. Besides collections of morphologically and syntactically-annotated data like Almaty Corpus of Kazakh Language [^3] http://web-corpora.net/KazakhCorpus/search/?interface_language=en and UD treebank Tyers and Washington (2015); Makazhanov et al. (2015), there exist also more task-specific text corpora.
@@ 4.3 Uzbek Language Datasets
Other Datasets. Some additional datasets, including multimodal and/or general unannotated data also exist. For example, one of the multimodal datasets is the open-source Uzbek Speech Corpus Musaev et al. (2021). On the other hand, there is the Uzbek corpus [^4] https://uzbekcorpus.uz/, which includes a large collection of educational, scientific, official, and artistic texts together with a morphological database and dictionaries, and an Uzbek Community corpus [^5] https://corpora.uni-leipzig.de/en?corpusId=uzb_community_2017, collected by Leipzig University and containing over 660 thousand sentences of community data only.
@@ 4.4 Kyrgyz Language Datasets
Unannotated Datasets and Corpora Collections. One of the largest open-source datasets for Kyrgyz is the Manas-UdS corpus [^6] https://fedora.clarin-d.uni-saarland.de/kyrgyz/index.html of over 84 literary texts in 5 genres, marked with lemmas and parts of speech Kasieva et al. (2020). For most other datasets, linking them directly to publications is not possible; however, some of them are available on Github. One of those is the KyrgyzNews dataset [^7] https://github.com/Akyl-AI/Kyrgyz_News_Corpus with over 250 thousand scraped news. Leipzig University has also compiled a corpus of over 3 million sentences from publicly available web sources. [^8] https://corpora.uni-leipzig.de/de?corpusId=kir_news_2020
Task-specific datasets. The few available task-specific datasets also can only be found on Github. One such example, for the task of NER, is the NER dataset [^9] https://github.com/Akyl-AI/KyrgyzNER that is currently under development by the researchers from Kyrgyz State Technical University (KSTU). Another example is the hand-written letters dataset [^10] https://github.com/Akyl-AI/kyrgyz_MNIST (Kyrgyz MNIST equivalent) also available on Github.
@@ 4.5 Turkmen Language Datasets
The data landscape of Turkmen language is even more scarce. Besides the Leipzig University’s corpora [^11] https://corpora.wortschatz-leipzig.de/en?corpusId=tuk-tm_web_2019 of over 270 thousand sentences of web-scraped data, other resources are practically non-existent. With only a few dictionaries and poetry and literature collections [^12] https://github.com/tmLang-NLP/datasets collected by enthusiasts and available in an open-source fashion on Github, Turkmen language falls largely behind every other Turkic language of Central Asia in terms of data availability.
@@ 4.6 Web-Scraped Datasets
All the above-mentioned languages are also represented in multilingual datasets predominantly scraped from the web. In particular, all the languages are available in CC100 Wenzek et al. (2020), WikiAnn Pan et al. (2017), and OSCAR Ortiz Suárez et al. (2020) datasets. Most of other web-scraped datasets present online are either subsets of the bigger multilingual datasets (like CC100 or OSCAR), or are scrapes from newspapers and websites available on the Internet, with many of them not being open-source (for example, the kkWaC dataset [^13] https://www.sketchengine.eu/kkwac-kazakh-corpus/ with over 139 million Kazakh words). However, as noted by Kreutzer et al. (2022), such datasets should be used with caution, as the quality of low-resource language data in them may be significantly lower. With only a few hundred labeled instances, even a 5-10 percent rate of mislabeled or grammatically incorrect sentences can substantially impact model performance in these languages.
@@ 4.8 Parallel Corpora
Additionally, we would like to comment on the availability of parallel corpora for the languages studied. While OPUS provides a comprehensive overview of parallel data available Tiedemann and Thottingal (2020) as well as several models Tiedemann et al. (2023), only a few less resourceful parallel pairs datasets exist, including the already mentioned Kazakh-Russian sign language corpora, Uzbek-Kazakh (Allaberdiev et al., 2024) and Kazakh-Russian (Kozhirbayev and Islamgozhayev, 2023) parallel corpora for MT and ASR.
@@ 4.9 Classifying Languages by Data Availability
The resources available for each language are far from abundant. If one were to categorize them according to the classification suggested by Joshi et al. (2020), Kazakh would most likely fit into “The Rising Star” category, with its substantial presence in the web and good variety of multimodal datasets. However, the research on this language is pushed back by a lack of labeled data for downstream tasks. Uzbek, given its recent developments and greater variability in terms of the linguistic resources available, would be categorized as “The Hopeful,” given that in the next years the efforts for collecting the datasets will not fade. Kyrgyz and Turkmen, unfortunately, not being sufficiently backed up by streamlined research efforts, would be classified as “The Scraping-Bys,” with the future of their data collection processes yet unclear.
@@ 5.2 Internet Access
Limited Internet access is the second reason for data scarcity in the region. As stated before, while a substantial amount of data has been gathered by NLP researchers from the Internet based sources, only 38 and 21 percent of users in Kyrgyzstan and Turkmenistan respectively have access to the global net. [^14] https://blogs.worldbank.org/en/europeandcentralasia/how-central-asia-can-ensure-it-doesnt-miss-out-digital-future While the situation is somewhat better in Uzbekistan (with 55 percent of population being able to access the medium) and significantly better in Kazakhstan (around 79 percent of population), limited connectivity hinders users’ abilities to contribute to open-source encyclopedias, blogs, news sites, and more. Additionally, the lack of digitalized resources, such as electronic books, journals, audio transcripts, and video recordings, restricts information sharing within the region.
@@ 8.1 Kazakh Language Technologies
Linguistic Analysis and Rule-Based Systems. A variety of tools have been developed for linguistic analysis and normalization of texts in Kazakh Yessenbayev et al. (2020). Early on, rule-based translation systems and morphological analyzers have also been developed Forcada and Tyers (2016). This research paved the way for the first advances in the most prominent areas of NLP, like machine translation and ASR, that are well represented in Kazakh language.
@@ 8.2 Uzbek Language Technologies
ASR. While Uzbek does not enjoy the same variety of machine translation tools as Kazakh, presence of automatic speech recognition technology is somewhat comparable with the most recent work contributed by Musaev et al. (2021), with their model performing at the level of 14.3% word error rate.
Fine-Tuning Existing Models. Uzbek also enjoys a better variety of pre-trained and fine-tuned models, including UzBERT Mansurov and Mansurov (2021), capable of outperforming mBERT by leveraging at least 11 times more language specific data; UzRoberta, pre-trained on roughly 2 million news articles Davronov and Adilova (2024); BERTBek, a model improving on UzBERT by training on Latin script Kuriyozov et al. (2024), and a compact and fine-tuned variation of a mT5 model [^15] https://ijdt.uz/index.php/ijdt/article/view/104. In general, with Uzbek still catching up on data availability, there is potential for the language to soon enjoy a better variety of NLP technologies.
@@ 8.3 Kyrgyz and Turkmen Languages Technology
Unfortunately, the situation for Kyrgyz and Turkmen looks less promising. With little to no work in terms of machine translation or automatic speech recognition, Kyrgyz enjoys little variety of technology for text classification Alekseev et al. (2023) and NER fine-tuned on WikiAnn data [^16] https://huggingface.co/murat/kyrgyz_language_NER. Turkmen offers even less language-specific technology, since it is has been mainly studied within the scope of comparative studies of Turkic languages. For example, some work on machine translation evaluation was done in the effort to build the infrastructure for Turkic languages, but nowadays this approach seems outdated (Alkim and Çebi, 2019). Overall, neither fine-tuning or pre-training of already available models and architectures have been researched for the majority of basic NLP tasks both in Kyrgyz and Turkmen.
@@ 9 Future Work Areas
Additionally, given the incorporation of LLMs in the field, potential of using them for data augmentation or annotation can be another important direction of research in comparison with more expensive methods of human data collection and annotation.
@@ 11 Limitations
Appendix A Search Strategy
A.1 General Approach
Firstly, we queried for papers using generic search terms and engines specified in A.3. Secondly, we focused on exploring popular CL and NLP venues, specific institutes’ resources (e.g. ISSAI’s website) as well as lists of references of works found in the first step of the process. Finally, we queried for pre-prints and unpublished works and datasets on arXiv, Github, Kaggle, and HuggingFace.
A.2 Venues and Repositories
The primary focus of our search was popular CL and NLP venues, including ACL, EACL, CoNLL, EMNLP, and WMT. We also explored non-ACL events and proceedings, including COLING and LREC.
For speech-related technology and datasets we researched works published at IEEE venues, including Interspeech, ICASSP, and SLT.
Additionally, we browsed pages of the universities and institutes that have been known for their contributions to the field, including ISSAI, Leipzig University, and Saarland University.
For pre-prints, works in progress, models, and datasets, we explored Github, Kaggle, and HuggingFace as well as the "Computation and Language" section on arXiv.
A.3 Search Engines and Query Details
Search engines used include Google Scholar, Semantic Scholar, and ResearchGate. We used both broad and specific query terms to search for publications on both resources and technologies. In terms of broad queries, we used the format of "[language] [topic]", for example, "Kyrgyz NLP" or "Kazakh speech". For more technology-specific searches we adopted queries of the form of "[language] [technology name/data]", for example, "Kazakh NER" or "Uzbek speech corpus".
The search cutoff date is set to November 1, 2024.
Appendix B Datasets